# Discrete Diffusion in Large Language and Multimodal Models: A Survey

### v3

@@ V.D.3 Reward Guidance
V.D.4 Energy-Based Diffusion
@@
VI Applications
VI.A Text Generation and Style Control
@@
[54] uses a diffusion-based language model for fine-grained text style transfer, trained on the StylePTB benchmark and achieving incredible results even with limited data and without external knowledge. [55] poses adopting diffusion language models for text embeddings, motivated by their inherent bidirectional architecture and recent success in matching or surpassing LLMs especially on reasoning tasks. Building on the idea of leveraging diffusion for controllable generation, [56] propose the Segment-Level Diffusion (SLD) framework, which segments long outputs and decodes them sequentially with an autoregressive decoder; as a result, diffusion predictions are simplified while coherence and scalability improve, and experiments confirm SLD’s fluency and contextual superiority over both diffusion and autoregressive baselines. For the social media applications, [57] introduces DiffusPoll, a model that generates diversified, high-quality poll options from user comments by combining a task-specific masking strategy with attribute-tag guidance; consequently DiffusPoll not only maintains quality and diversity but also better captures minority viewpoints. In the creative-writing domain, [58] present PoetryDiffusion, which jointly enforces semantic meaning and metrical form. By generating entire lines in a single pass and integrating a novel metrical-control module, PoetryDiffusion flexibly adjusts rhyme and structure, and experiments show it produces poems that meet both semantic and metrical requirements with high quality.
VI.B Text Editing and Summarization
EdiText [59] is a controllable text-editing framework that operates at both coarse and fine levels to achieve target attributes; specifically, it fuses an SDEdit-style module for large-scale edits with a novel self-conditioning mechanism for subtle refinements, thereby enabling substantial stylistic changes while preserving meaning in tasks such as toxicity and sentiment control. Moving from editing to long-form generation, [60] designs a discrete diffusion model for efficient abstractive summarization, where a semantic-aware noising schedule, together with an adapted CrossMamba backbone, allows transformers to process long sequences under diffusion and yields strong results on Gigaword and CNN/DailyMail. Complementing summarization with unsupervised exploration of document themes, DiffETM [61] injects diffusion into the Embedded Topic Model so that the document–topic distribution is sampled via a more realistic stochastic process—consequently, the model better fits true topic structures while keeping optimization tractable, as confirmed on two standard corpora. Furthermore, focusing on domain-specific summarization, TermDiffuSum [62] leverages term-aware attention during diffusion and a re-ranking loss to highlight legally salient sentences, thereby surpassing previous diffusion-based extractive methods on legal-document benchmarks.
VI.C Sentiment Analysis and Data Augmentation
@@
The $\textrm{CDA}^{2}$ framework [63] employs counterfactual diffusion augmentation to improve cross-domain adaptation in low-resource sentiment analysis. It generates counterfactual target samples by substituting domain-relevant words in source data and guiding the diffusion model to produce high-quality target domain examples. Experiments show that $\textrm{CDA}^{2}$ generates higher-quality target samples and achieves state-of-the-art performance in cross-domain sentiment tasks. DiffusionCLS [64] examines the use of a diffusion language model for data augmentation in low-resource sentiment classification. It generates pseudo-samples by reconstructing strongly sentiment-related tokens, thereby increasing data diversity while preserving label-consistent information. This approach balances consistency and diversity without introducing excessive noise. Pushing diffusion further into structured sentiment tasks, the Grid Noise Diffusion Pinpoint Network (GDP) [65] introduces a novel diffusion-based and template-guided approach for Aspect Sentiment Quad Prediction. It includes three modules: Diffusion Vague Learning to enhance learning under uncertainty, Consistency Likelihood Learning to capture commonalities of sentiment elements and mitigate noise, and a GDP-FOR generation template for more natural outputs. Extensive experiments demonstrate the remarkable effectiveness of GDP on ASQP tasks. [106] addresses the layout sticking issue in discrete diffusion layout generation, where poorly placed elements become hard to fix. It proposes Layout-Corrector, a simple module that scores elements’ layout quality, resets poorly positioned tokens, and guides refined placement conditioned on high-quality elements. Combined with discrete diffusion models, Layout-Corrector consistently enhances layout quality, flexibility, and sampling efficiency.
VI.D Knowledge and Reasoning
@@
Diffusion-of-Thought (DoT) [66] firstly integrates chain-of-thought reasoning into dLLMs to enhance reasoning capabilities. Unlike traditional left-to-right autoregressive reasoning, DoT allows reasoning steps to diffuse over multiple steps, offering a flexible trade-off between compute and reasoning performance. In addition, DiffuCOMET [67] develops a series of models that leverage diffusion to infer contextually relevant commonsense knowledge from narratives. The method progressively refines a commonsense fact representation anchored to the input context through multiple diffusion steps, producing inferences that are both contextually appropriate and diverse. On benchmarks like ComFact and WebNLG+, DiffuCOMET achieves a better trade-off between commonsense diversity, contextual relevance, and alignment to known facts. DPCL-Diff [68] combines graph node diffusion with dual-domain periodic contrastive learning for temporal knowledge graph reasoning. Its Graph Node Diffusion (GNDiff) model introduces noise into sparsely related historical events to generate new event data that better matches the true distribution, enhancing the ability to reason about future events; the dual-domain learning maps periodic event entities to a Poincaré space and non-periodic entities to a Euclidean space to improve representation power. Most recently, the d1 framework [69] adapts pretrained dLLMs into reasoning models via a combination of supervised fine-tuning and reinforcement learning. It introduces a novel critic-free policy-gradient algorithm called diffu-GRPO and employs masked SFT to distill reasoning knowledge directly from existing datasets, thereby enhancing the model’s reasoning abilities. [49] provides an insight into DLLM reasoning: the difficulty of decoding individual tokens in a response varies, and it is not necessarily the case that tokens on the left are easier to decode than those on the right. Therefore, autoregressive (AR) reasoning may not be the optimal approach. Diffusion-based LLMs demonstrate clear advantages over AR models in tasks involving planning, reverse reasoning, and global context understanding. [49] presents examples such as Sudoku and the 24 Game to illustrate these points. NeSyDM [107] introduces a discrete diffusion process in the symbolic space to overcome the limitations of traditional neurosymbolic systems, which often rely on the assumption of conditional independence between symbols. The framework first extracts symbolic representations from perceptual inputs using a neural network, and then performs a multi-step denoising diffusion process over these symbols. This allows the model to explicitly capture inter-symbol dependencies through the structure of the diffusion trajectory.
VI.E Vision and Multimodal
DiffVLA [70] introduces a vision-language-guided diffusion policy for autonomous-driving planning. Specifically, it employs a hybrid sparse–dense diffusion strategy, guided by a vision-language model, thereby improving both efficiency and action diversity in multi-modal driving scenarios. In a image-enhancement context, UDAN-CLIP [71] proposes an image-to-image diffusion framework for underwater enhancement that jointly integrates a vision-language model and a spatial-attention module. During pre-training, the model leverages synthetic underwater data and a CLIP-guided loss to retain in-air natural priors while correcting local degradations such as haze and low contrast. Extending diffusion policies to robotics, [72] present a framework VPDD that first pre-trains on large-scale actionless human videos and then transfers the learned discrete diffusion policy to robot-control tasks. Here, both human and robot videos are encoded into unified video tokens; a discrete diffusion model is trained generatively on the human corpus before being fine-tuned on a small, action-labeled robot set. Turning to motion synthesis, M2D2M [73] employs a discrete diffusion model to generate continuous human-motion sequences from textual descriptions of multiple actions. By adapting dynamic transition probabilities to token-level similarities and adopting a two-phase sampling strategy—independent followed by joint denoising—it produces long-term, smooth, and contextually coherent motions, ultimately outperforming state-of-the-art multi-motion baselines. Finally, AR-Diffusion [74] introduces a novel architecture that blends autoregressive and diffusion techniques for flexible, asynchronous video generation. It gradually corrupts video frames during both training and inference to reduce phase discrepancies; meanwhile, a non-decreasing length-control mechanism—borrowed from autoregressive generation—enables consistent handling of variable-length sequences, thereby enhancing temporal coherence. Besides, the understanding tasks, discrete diffusion is also largely used in the vision generation tasks [47, 103].
VI.F Biological and Drug Discovery
@@
Several recent studies leverage dLLMs and dMLLMs to advance molecular editing, protein engineering, and drug discovery. MolEditRL [75] is a molecular-editing framework that combines a discrete graph-diffusion model with reinforcement learning to optimize molecular properties while preserving structural similarity. It operates in two stages: first, a conditional diffusion model reconstructs a target molecule from the source structure and textual instructions; second, reinforcement learning fine-tunes the editing actions, thereby further improving property alignment and structural conservation. Building on the idea of diffusion for biomacromolecules, CFP-Gen [76] adopts a diffusion language model for combinatorial functional protein generation, thus enabling de novo design under simultaneous functional, sequence, and structural constraints. It introduces two key modules—Annotation-Guided Feature Modulation (AGFM) and Residue-Controlled Functional Encoding (RCFE)—that dynamically adjust protein features according to functional annotations and precisely control residue-level interactions. As a result, CFP-Gen can create novel proteins whose functionality rivals that of natural counterparts, and it attains a high success rate in designing multifunctional proteins. In a related vein, TransDLM [77] proposes a text-guided, multi-property molecular-optimization method that leverages a diffusion language model. By encoding molecules with standardized chemical nomenclature and embedding property requirements directly into textual descriptions, TransDLM implicitly enforces multiple objectives and thereby reduces error propagation during optimization. Further generalizing diffusion to drug discovery, GenMol [78] presents a single, discrete diffusion model that serves as a versatile generator across diverse pharmaceutical tasks. It produces Sequential Attachment-based Fragment Embedding (SAFE) sequences via non-autoregressive bidirectional decoding, thus avoiding token-order constraints and boosting sampling efficiency. GenMol also treats molecular fragments as basic building blocks and introduces a fragment-remasking strategy to refine candidate structures. Addressing sequence–structure co-design, DPLM-2 [79] is a multimodal protein language model capable of understanding and generating both protein sequences and their three-dimensional structures. It converts 3-D coordinates into discrete tokens through quantization, then jointly trains on sequence and structure data, thereby capturing complex sequence-structure relationships and improving tasks that demand structure-conditioned sequence generation. PepTune [80] targets therapeutic-peptide design with a multi-objective, discrete diffusion framework built on a masked language-model backbone. It introduces bond-dependent masking and a Monte-Carlo Tree Guidance algorithm that balances exploration and exploitation during inference, iteratively refining peptides toward multiple goals. Consequently, PepTune generates chemically diverse peptides that are simultaneously optimized for properties such as binding affinity, membrane permeability, solubility, and hemolysis.
VII Future Directions
In this survey, we have reviewed the recent progress of dLLMs and dMLLMs, and summerized the key mathematics, base models, and training/inference techniques for understanding and utilizing this type of model. We mainly focus on the large-sized models while shortly introducing the contents of early small-sized models (e.g., D3PM, Plaid, SEDD, etc.) that have been well explored and covered in the existing literature. Next, we introduce the challenges and future directions for dLLMs and dMLLMs in the following aspects.
VII.A Training and Infrastructure

### v4

@@ VI.D.3 Reward Guidance
VI.D.4 Energy-Based Diffusion
@@
VI.E Sampling Techniques
VI.E.1 Temporal Self-Consistency
[74] identifies a phenomenon termed temporal oscillation, where correct intermediate predictions emerge during the iterative decoding steps but are later overwritten by incorrect outputs in subsequent steps.
To mitigate the performance degradation caused by temporal oscillation, rather than relying exclusively on the final denoising step, [74] utilizes the predictions at all timesteps and predicts through a weighted voting scheme:
$\text{answer}^{\ast}=\arg\max_{a}\sum_{t=1}^{T}f(t)\cdot\mathbf{1}\big{(}\text{semantic meaning}(x^{t}_{0})=a\big{)},$ (118)
where $x^{t}_{0}$ denotes the prediction at time step $t$, and $\text{semantic meaning}(x^{t}_{0})$ represents the extracted answer from $x^{t}_{0}$, $\mathbf{1}(\cdot)$ is the indicator function and $f(t)$ is a weighting function (e.g., exponential decay) assigning more importance to later steps.
[74] also introduces Temporal Semantic Entropy (TSE), a measure of semantic uncertainty across intermediate predictions. The computation os TSE is based on the semantic clustering of historical answers. This metric can be incorporated into the reinforcement learning advantage function, thereby mitigating the issue of temporal oscillation in the model.
VI.E.2 Particle Gibbs Sampling for Discrete Diffusion Models
Instead of sampling only once to obtain a single trajectory, [73] introduces a reference trajectory that is iteratively refined across multiple rounds of sampling.
The method begins by drawing an initial reference trajectory $x^{\prime}_{0:T}$ using Sequential Monte Carlo (SMC) with a single particle. At each subsequent iteration, a set of $k$ particles is generated, where one particle is deterministically fixed to the reference trajectory while the remaining k1k-1 particles are defined as the candidate particles and are initialized randomly. At each step $t$, candidate particles are propagated according to the model’s denoising distribution,
$\bar{x}^{(i)}_{t-1}\sim p_{\theta}(x_{t-1}\mid c,x^{(i)}_{t})$, (119)
with the reference particle $\bar{x}^{(k)}_{t-1}$ kept fixed. The importance weight are for each candidate particles at time $t$ is defined based on a reward function $r$:
$w^{(i)}_{t-1}=\exp\left(\frac{r(c,\bar{x}^{(i)}_{t-1})-r(c,x^{(i)}_{t})}{\beta}\right),$ (120)
and normalized before resampling. After resampling, the reference trajectory is updated by selecting one trajectory from the set with probability proportional to its importance weight. Iterating this procedure $m$ times yields a refined trajectory distribution that converges to the reward-weighted posterior.
VI.E.3 Early Stopping
[27] finds that diffusion models can often generate the correct answer before completing the entire decoding process. Therefore, discrete diffusion can leverage early stopping to improve decoding efficiency. Both [27] and [28] employ either predefined response templates or in-place prompts to divide the response space into two parts: the final answer region and the reasoning region. Once the confidence in the final answer region reaches a sufficient level, the subsequent decoding is terminated.
VI.F Context Length Extension
A notable phenomenon observed in diffusion-based large language models is that, with bidriectional attention, their robustness in handling extended contexts. Unlike auto-regressive models that suffer catastrophic degradation when the context length exceeds the pretraining window, diffusion LLMs can still reliably retrieve information from the most recent segment of the input, even at depths far beyond the training horizon.
To further extend the context window, LongLLaDA [75] applies the NTK-aware scaling method to RoPE within diffusion LLMs. Following [152], the critical dimension $d_{\text{extra}}$ is defined as:
$d_{\text{extra}}=2\cdot\left\lceil\frac{d}{2}\cdot\log_{\beta_{0}}\left(\frac{T_{\text{train}}}{2\pi}\right)\right\rceil,$ (121)
where $d$ is the hidden dimension, $\beta_{0}$ is the rotary base, and $T_{\text{train}}$ is the pretraining context length. Given a target context length $t$, the NTK scaling factor $\lambda$ is chosen as:
$\lambda=\beta_{0}^{-1}\cdot\left(\frac{t}{2\pi}\right)^{d/d_{\text{extra}}}.$ (122)
VI.G Sparse Computation
The forward computation of diffusion-based large language models exhibits inherent sparsity, allowing inference efficiency to be improved by eliminating part of the computations.
Under the block-wise decoding setup, many techniques have been explored. Sparse-dLLM [76] indicates that it is unnecessary to cache and reuse all key–value pairs of tokens outside the current block. By applying an average attention score between non-current tokens and tokens in the current block, a subset of tokens can be filtered out, thereby reducing the memory consumption of the KV cache and improving inference throughput. DPad [77] demonstrates that tokens in subsequent blocks do not need to attend to all tokens in the current block. Instead, a random dropout can be applied, with a higher dropout rate assigned to tokens that are farther from the current block. Pipelined Parallel decoding [78] adopts a “soft” block-wise decoding strategy. Instead of waiting for the current block to be fully decoded before moving to the next, it begins decoding tokens in the subsequent block once the proportion of decoded tokens in the last active block exceeds a predefined threshold. This gradual inclusion of additional blocks reduces computation on the right-side tokens and thereby improves efficiency.
VI.H Response Length Control
Adaptive control of response length is an essential capability in discrete diffusion models, and it can be achieved through either training-based or training-free approaches. In training-based methods, DreamOn [51] enables dynamic adjustment of token counts by introducing two special tokens, expand and delete. Similarly, FlexMDM [39] employs an auxiliary network to predict how many tokens should be inserted before each existing token, thereby allowing the response length to grow dynamically. In contrast, the following section introduces training-free techniques for adaptive response length control.
DAEDAL [79] introduces a two-stage adaptive mechanism for response length control. It uses the confidence in predicting the End-of-Sequence (EOS) token as a signal of length sufficiency. In the first stage, the model begins with a short initial sequence length $L_{\text{init}}$. A preliminary length estimation loop iteratively evaluates EOS confidence. Whenever the confidence falls below $\tau_{\text{eos}}$, the sequence is expanded by appending $E_{\text{factor}}$ additional mask tokens until either $\tau_{\text{eos}}$ is satisfied or a maximum length $L_{\max}$ is reached. In the second stage, the model proceeds with standard iterative denoising. However, in addition to filling high-confidence tokens, DAEDAL introduces an expansion mechanism triggered by low-confidence predictions. Let $p_{\theta}(x_{t}[i])$ denote the confidence at position $i$. If $p_{\theta}(x_{t}[i])<\tau_{\text{low}}$, then the position with lowest confidence will serve as an expansion point and be replaced by a block of $E_{\text{factor}}$ mask tokens.
VII Privacy and Safety
Diffusion-based large language models (dLLMs) introduce unique safety vulnerabilities stemming from their bidirectional context modeling and parallel decoding mechanisms. These features, while enabling efficient infilling and interactive generation, also weaken the defenses that are effective in autoregressive (AR) LLMs.
Diffusion-based LLMs Jailbreak Attack (DIJA) [80] and PArallel Decoding jailbreak (PAD)[81] reformulates conventional jailbreak prompts into an interleaved mask-text format, compelling the model to generate unsafe outputs while maintaining contextual consistency. Formally, let $a=(a_{1},\ldots,a_{R})$ denote a vanilla jailbreak prompt, and let $m=([MASK],\ldots,[MASK])_{Q}$ represent $Q$ consecutive mask tokens. DIJA and PAD construct an interleaved prompt
$p_{i}=a\oplus(m\otimes w),$ (123)
where $\oplus$ denotes concatenation, $\otimes$ denotes interleaving, and $w$ represents benign separator text, such as, “Step 1” and ”Step 2”. Importantly, the hazardous intent contained in $a$ is preserved, while critical instructions are forced into masked positions.
[81] shows that dLLMs are more vulnerable to manipulation at the middle of the response than at the initial tokens and previous otimization-based jailbreak methods are effective at manipulating initial tokens but largely fail to optimize middle tokens. Thus, [81] proposes the Middle tOken Safety Alignment (MOSA), a reinforcement learning alignment strategy, which requires the model’s generated middle tokens to align with a set of predefined safe tokens. Experimental evidence shows that optimizing alignment in the middle portion of the sequence is more effective than optimizing alignment at the beginning.
VIII Quantization
[83] presents a systematic study of post-training quantization (PTQ) on dLLMs, benchmarking mainstream quantization methods across multiple models. Their analysis highlights severe activation outliers in dLLMs, sensitivity differences across tasks, and the relative robustness of instruction-tuned variants compared to base models.
[84] also analyzes why conventional PTQ methods degrade severely on dLLMs, identifying three key challenges: (1) quantization errors accumulate across iterations, (2) distinct token distributions across decoding steps, and (3) significant disparities in feature distributions across both token and channel dimensions. To address these, they introduce DLLMQuant, a framework featuring Temporal-Mask Adaptive Sampling (TMAS), Interaction-Aware Activation Quantization (IA-AQ), and Certainty-Guided Quantization (CGQ).
TMAS addresses the calibration challenge in dLLMs by accounting for variations across time steps and masking ratios. Specifically, it divides the iterative generation into blocks and selects calibration inputs at specific time intervals, ensuring coverage of diverse mask ratios across all timesteps.
Quantization errors $L(x_{t})$ at time step $t$ in dLLMs accumulate geometrically across denoising steps. Formally, the error propagation can be expressed as:
$L(x_{t})$ $=x_{t}-\text{Deq}\!\left(Q\!\left(x_{t}+L(x_{t+1})\right)\right)$ (124)
$=Q_{\text{Model}}(x_{t+1})-Q_{\text{Model}}\!\left(\text{Deq}\!\left(Q(x_{t+1})\right)\right)$
where $Q(\cdot)$ is the quantization process, Deq is the dequantization operation, and $Q_{\text{Model}}$ denotes the quantized model.
A key source of error is the matrix multiplication between softmax outputs and the value matrix in attention. IA-AQ redefines the quantization loss for value matrix $V$ as:
$L(s)=\left\|\Big{(}\lfloor\tfrac{V-z}{s}\rfloor-V\Big{)}\cdot\text{Deq}(O_{\text{softmax}})\right\|^{2}_{F},$ (125)
where $O_{\text{softmax}}$ is the softmax output, $z$ is zero-point, and $s$ is the scale factor. The optimal scaling factor is chosen by:
$s=\arg\min_{\alpha\in\{1.0,0.8\}}L(\alpha\odot\hat{s}),\quad\hat{s}=\frac{V_{\max}-V_{\min}}{Q_{\max}-Q_{\min}}.$ (126)
In addition, not all tokens equally affect subsequent iterations. Errors on unmasked or low-confidence tokens are small, while masked tokens with high confidence dominate the next step. To account for this, CGQ incorporates token certainty into the Hessian used for weight quantization:
$H=\big{(}X\odot(1[X_{t}=M]+\sqrt{sc_{t}})\big{)}\cdot\big{(}X\odot(1[X_{t}=M]+\sqrt{sc_{t}})\big{)}^{\top},$ (127)
where $1[X_{t}=M]$ indicates a custom weighted indicator function, and $sc_{t}$ is the final confidence score to each token in model output. This certainty-weighted Hessian prioritizes minimizing quantization error on critical masked tokens.
IX Applications
IX.A Language Related Applications
@@
[85] introduce a diffusion-based language model for fine-grained text style transfer. [86] propose adopting diffusion models for text embeddings, motivated by their bidirectional architecture and recent successes in matching or surpassing LLMs, particularly on reasoning tasks. Extending diffusion for controllable generation, [87] present the Segment-Level Diffusion (SLD) framework, which divides long outputs into segments and decodes them sequentially with an autoregressive decoder. For social media applications, [88] introduce DiffusPoll, a model that generates diverse, high-quality poll options from user comments by combining a task-specific masking strategy with attribute-tag guidance. In creative writing, [89] propose PoetryDiffusion, which jointly enforces semantic meaning and metrical form. Similarly, [90] employ masked diffusion language models for controllable style transfer, introducing inference-time verifier guidance that improves stylistic alignment without retraining.
EdiText [91] extends controllable generation to text editing, operating at both coarse and fine levels to achieve target attributes. Moving from editing to long-form generation, [92] design a discrete diffusion model for abstractive summarization. Complementing this direction, DiffETM [93] inject diffusion into the embedded topic model, enabling document–topic distributions to be sampled through a more realistic stochastic process. Focusing on domain-specific summarization, TermDiffuSum [94] integrates term-aware attention and a re-ranking loss during diffusion, effectively emphasizing legally salient sentences.
@@
The $\textrm{CDA}^{2}$ framework [95] employs counterfactual diffusion augmentation to improve cross-domain adaptation in low-resource sentiment analysis. DiffusionCLS [96] examines the use of a diffusion language model for data augmentation in low-resource sentiment classification. It generates pseudo-samples by reconstructing strongly sentiment-related tokens, thereby increasing data diversity while preserving label-consistent information. The Grid Noise Diffusion Pinpoint Network (GDP) [97] introduces a novel diffusion-based and template-guided approach for Aspect Sentiment Quad Prediction. [98] addresses the layout sticking issue in discrete diffusion layout generation, where poorly placed elements become hard to fix. It proposes Layout-Corrector, a simple module that scores elements’ layout quality, resets poorly positioned tokens, and guides refined placement conditioned on high-quality elements.
IX.B Reasoning
@@
Diffusion-of-Thought (DoT) [99] firstly integrates chain-of-thought reasoning into dLLMs to enhance reasoning capabilities. DoT allows reasoning steps to diffuse over multiple steps, offering a flexible trade-off between compute and reasoning performance. In addition, DiffuCOMET [100] develops a series of models that leverage diffusion to infer contextually relevant commonsense knowledge from narratives. DPCL-Diff [101] combines graph node diffusion with dual-domain periodic contrastive learning for temporal knowledge graph reasoning. Most recently, the d1 framework [102] adapts pretrained dLLMs into reasoning models via a combination of supervised fine-tuning and reinforcement learning. It introduces a novel critic-free policy-gradient algorithm called diffu-GRPO and employs masked SFT to distill reasoning knowledge directly from existing datasets, thereby enhancing the model’s reasoning abilities. One recent work [60] provides an insight into dLLM reasoning: the difficulty of decoding individual tokens in a response varies, and it is not necessarily the case that tokens on the left are easier to decode than those on the right. NeSyDM [103] introduces a discrete diffusion process in the symbolic space. The framework first extracts symbolic representations from perceptual inputs using a neural network, and then performs a multi-step denoising diffusion process over these symbols.
IX.C Vision and Multimodal
UDAN-CLIP [104] proposes an image-to-image diffusion framework for underwater enhancement. Turning to motion synthesis, M2D2M [105] employs a discrete diffusion model to generate continuous human-motion sequences from textual descriptions of multiple actions. AR-Diffusion [106] introduces a novel architecture that blends autoregressive and diffusion techniques for flexible, asynchronous video generation. Besides the understanding tasks, discrete diffusion is also largely used in the vision generation tasks [57, 144].
IX.D Robotics and Autonomous Driving
DiffVLA [107] introduces a vision-language-guided diffusion policy for autonomous-driving planning. ViLaD [109] introduces a large vision–language diffusion framework for end-to-end autonomous driving, addressing the latency and unidirectional limitations of autoregressive VLM-based decision models. Meanwhile, Discrete Diffusion VLA [108] unifies vision, language, and action decoding in a single transformer. It discretizes actions into tokens and employs iterative re-masking for adaptive, parallel action generation. Extending diffusion policies to robotics, [110] present a framework VPDD that first pre-trains on large-scale actionless human videos and then transfers the learned discrete diffusion policy to robot-control tasks. Discrete-Guided Diffusion (DGD) [111] integrates discrete multi-agent path finding (MAPF) with diffusion models to address scalable and safe multi-robot motion planning. By using MAPF solutions as structured priors and convex decomposition for local planning, DGD enables parallel denoising while ensuring feasibility through lightweight constraint projection.
IX.E Graph and Structured Predictions
LO-ARM [112] introduces a learning-order autoregressive model that dynamically adapts generation order for molecular graphs, improving flexibility and validity. ReDiSC [113] proposes a reparameterized discrete masked diffusion model for node classification. By modeling structured label dependencies with variational EM and GNN integration, ReDiSC achieves scalable and interpretable predictions on large graphs. Scaffold Diffusion [114] formulates sparse multi-category voxels as discrete token sequences and applies a dLLM for generation. With 3D sinusoidal positional encoding and occupancy-map constraints, it produces functionally consistent voxel structures. This work demonstrates the potential of dLLMs for 3D sparse structure generation, though it remains limited by sequence length and may benefit from hierarchical extensions.
IX.F Biological and Drug Discovery
@@
MolEditRL [115] is a molecular-editing framework that combines a discrete graph-diffusion model with reinforcement learning to optimize molecular properties while preserving structural similarity. Building on the idea of diffusion for biomacromolecules, CFP-Gen [116] adopts a diffusion language model for combinatorial functional protein generation It introduces two key modules—Annotation-Guided Feature Modulation (AGFM) and Residue-Controlled Functional Encoding (RCFE)—that dynamically adjust protein features according to functional annotations and precisely control residue-level interactions. TransDLM [117] proposes a text-guided, multi-property molecular-optimization method that leverages a diffusion language model. TransDLM implicitly enforces multiple objectives and thereby reduces error propagation during optimization. Further generalizing diffusion to drug discovery, GenMol [118] presents a single, discrete diffusion model that serves as a versatile generator across diverse pharmaceutical tasks. Addressing sequence–structure co-design, DPLM-2 [119] is a multimodal protein language model capable of understanding and generating both protein sequences and their three-dimensional structures. PepTune [120] targets therapeutic-peptide design with a multi-objective, discrete diffusion framework built on a masked language-model backbone. It introduces bond-dependent masking and a Monte-Carlo Tree Guidance algorithm that balances exploration and exploitation during inference, iteratively refining peptides toward multiple goals. [121] propose CMCM-DLM, which integrates structure-control and property-control modules into a pretrained diffusion language model for molecules.
X Future Directions
X.A Training and Infrastructure